# Connectedness of level sets for non-degenerate integrable systems that extend complexity one torus actions

### v2

@@ 7. Proof of the main result
If there are no non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer, then by Theorem 1.6 $f^{-1}(\alpha,c)$ is connected for all $(\alpha,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$. So assume instead that $f^{-1}(\alpha,c)$ is connected for all $(\alpha,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$. Fix $\alpha\in{\mathfrak{t}}^{*}$. Assume without loss of generality that $\alpha=0$. Since $\Phi$ is proper, the reduced space ${M/\!/T}$ is compact and connected by Theorem 5.5.
If the reduced space ${M/\!/T}$ is a single point then every point in the orbit $\Phi^{-1}(0)$ is short, and so by Lemma 6.3, every non-degenerate singular point in $\Phi^{-1}(0)$ has purely elliptic type; in particular, it has no hyperbolic block.
Otherwise, by Proposition 7.1, ${M/\!/T}$ is a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function; moreover, given $p\in\Phi^{-1}(0)$, $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer. If $\Phi^{-1}(\beta)/T$ is simply connected for some $\beta\in\Phi(M)$, then $\Phi^{-1}(\alpha)/T$ is simply connected by [KT1, Lemma 5.7 and Corollary 9.7]. Therefore, there are no non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer by Lemma 7.3. Similarly, if $f^{-1}(\alpha,c)$ contains at most one orbit of non-degenerate singular points with a hyperbolic block for all $c\in{\mathbb{R}}$, then it contains no non-degenerate singular points with hyperbolic block and connected $T$-stabilizer by Lemma 7.4.
@@
Let $N$ be a connected manifold and let $h\colon N\to{\mathbb{R}}$ be a proper function such that $h(N)\subseteq[0,\infty)$. Assume that the restriction of $h$ to $h^{-1}(0,\infty)$ is a Morse-Bott function with no points of index $1$, and that 00 is an isolated critical value. Then $h^{-1}(0)$ is connected.
@@
Consider points $x$ and $y$ in $h^{-1}(0)$. Since $N$ is path-connected, there is a path $\gamma$ from $x$ to $y$. Hence, $x$ and $y$ are in the same component of $h^{-1}([0,b))$ for some $b\in{\mathbb{R}}$. If $C\subseteq h^{-1}((0,\infty))$ is a critical submanifold, then since the index of $C$ is not $1$, the restriction $H^{0}(h^{-1}([0,f(C)+\epsilon)))\to H^{0}(h^{-1}([0,f(C)-\epsilon)))$ is a surjection for sufficiently small $\epsilon>0$. In particular, if $x$ and $y$ are in the same component of $h^{-1}([0,h(C)+\epsilon))$, then they are also in the same component of $h^{-1}([0,h(C)-\epsilon))$. Moreover, since $h$ is proper and 00 is an isolated critical value, there are a finite number of such critical submanifolds in $h^{-1}([0,b))$. Hence, by induction, $x$ and $y$ are in the same component of $h^{-1}(0).$
@@
Fix $c\in{\mathbb{R}}$ and define $h\colon N\to{\mathbb{R}}$ by $h=(\overline{g}-c)^{2}$. On the set $\overline{g}^{-1}(c,\infty)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide and have the same index. On the set $\overline{g}^{-1}(-\infty,c)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide but the index and coindex of each critical manifold are exchanged. Finally, since $h$ is a proper Morse-Bott function there exists a neighborhood of $c$ that contains no other critical values, so 00 is an isolated critical value. Therefore, $h^{-1}(0)=\overline{g}^{-1}(c)$ is connected by Lemma 7.5. ∎

### v3

@@ 7. Proof of the main result
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. Fix $\beta\in{\mathfrak{t}}^{*}$ such that $\Phi^{-1}(\beta)/T$ contains more than one point; we may assume that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5. Hence, every point in M//T{M/\!/T} is tall, and so by Proposition 7.1, M//T{M/\!/T} can be given the structure of a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function with connected fibers; moreover, given $p\in\Phi^{-1}(0)$, the orbit $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer. Finally, since $T$ is connected, the fiber $\overline{g}^{-1}(c)$ is connected exactly if the fiber $f^{-1}(\beta,c)$ is connected for any $c\in{\mathbb{R}}$.
If $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M)$, then M//T{M/\!/T} is simply connected by [KT1, Lemma 5.7 and Corollary 9.7]. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemma 7.3.
Similarly, if each fiber of $f$ contains at most one $T$-orbit of non-degenerate singular points with a hyperbolic block, then each fiber of $\overline{g}$ contains at most one one critical point of index $1$. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemmas 7.3 and 7.4. ∎
@@
Let $N$ be a connected manifold and let $h\colon N\to{\mathbb{R}}$ be a proper function such that $h(N)\subseteq[0,\infty)$. Assume that the restriction of $h$ to $h^{-1}(0,\infty)$ is a Morse-Bott function with no points of index $1$, and that 0 is an isolated critical value. Then $h^{-1}(0)$ is connected.
@@
Consider points $x$ and $y$ in $h^{-1}(0)$. Since $N$ is path-connected, there is a path $\gamma$ from $x$ to $y$. Hence, $x$ and $y$ are in the same component of $h^{-1}([0,b))$ for some $b\in{\mathbb{R}}$. If $C\subseteq h^{-1}((0,\infty))$ is a critical submanifold, then since the index of $C$ is not $1$, the restriction $H^{0}(h^{-1}([0,f(C)+\epsilon)))\to H^{0}(h^{-1}([0,f(C)-\epsilon)))$ is a surjection for sufficiently small $\epsilon>0$. In particular, if $x$ and $y$ are in the same component of $h^{-1}([0,h(C)+\epsilon))$, then they are also in the same component of $h^{-1}([0,h(C)-\epsilon))$. Moreover, since $h$ is proper and 0 is an isolated critical value, there are a finite number of such critical submanifolds in $h^{-1}([0,b))$. Hence, by induction, $x$ and $y$ are in the same component of $h^{-1}(0).$
@@
Fix $c\in{\mathbb{R}}$ and define $h\colon N\to{\mathbb{R}}$ by $h=(\overline{g}-c)^{2}$. On the set $\overline{g}^{-1}(c,\infty)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide and have the same index. On the set $\overline{g}^{-1}(-\infty,c)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide but the index and coindex of each critical manifold are exchanged. Finally, since $h$ is a proper Morse-Bott function there exists a neighborhood of $c$ that contains no other critical values, so 0 is an isolated critical value. Therefore, $h^{-1}(0)=\overline{g}^{-1}(c)$ is connected by Lemma 7.5. ∎